OpenStudy (anonymous):

The Cost of producing x chairs is given by C(x) = .02x ^{2} + 40x + 400. Find C ^{1} (20)

OpenStudy (anonymous):

The Cost of producing x chairs is given by \[C(x) = .02x ^{2} + 40x + 400 Find C ^{1} (20)\]

OpenStudy (anonymous):

By \[C^{1}\]do you mean\[C \prime\]As in, the first derivative?

OpenStudy (anonymous):

would it be \[C ^{1} = 1/4x + 4 = C ^{1} (20) 4+5 = 9 ???\]

OpenStudy (anonymous):

\[Cprime1\]

OpenStudy (anonymous):

The second one

OpenStudy (anonymous):

So you are looking to find C prime? For the sake of moving this forward, that's what I will assume. In order to find the derivative of C(x), we would use the power rule. \[f(x)=x^{n}\]\[f \prime=nx^{n-1}\] You can treat each term in the polynomial as a different function. So you can apply the rule to each term and add them together. So apply the power rule to \[0.2x^{2}\]then apply the rule to \[40x\]and finally \[400\] Then add your results together.